İf the supply curve is q = 3 + 1.5p, what is the producer surplus if the price is 12?

1
Expert's answer
2021-10-31T18:30:05-0400

"Q=3+1.5p\\P=12\\Q=3+1.5(12)\\Q=21"

It will be a deficit of 75

Producer surplus

"=12\times 21-\int ^{21}_0[\frac{q^2-2q}{}-2q]^{21}_{0}\\=252-(147-42)\\=147"

producer surplus is 147 units